OpenStudy (anonymous):

WILL GIVE MEDAL!!!! A= (d1)(d2)/2 A=2a^2-8a d1=(a-4)

OpenStudy (amorfide):

what is the question?

OpenStudy (anonymous):

SOLVE FOR D2

OpenStudy (amorfide):

substitute your values in A= (d1)(d2)/2 A=2a^2-8a d1=(a-4) since you know what A is and what d1 is 2a^2-8a= (d1)(d2)/2 then sub in d1 2a^2-8a= (a-4)(d2)/2 rearrange to get d2, can you do that?

OpenStudy (anonymous):

i know the answer is 4a i just keep on getting something different

OpenStudy (amorfide):

4a^2 - 16a=(a-4)(d2) (4a^2-16a)/(a-4)=d2 4a(a-4)/(a-4)=d2 a-4 will cancel out to be 1 4a(1)=d2 d2=4a

OpenStudy (anonymous):

do you HAVE to factor (4a^2-16a)?

OpenStudy (amorfide):

well because we have (4a^2-16a)/(a-4) you would first try to factorise the top to see if it can cancel anything out to make it simpler, or you can use division to get the answer I would recommend factorising though since this question is basically asking for it

OpenStudy (anonymous):

Oh okay Thank you I've been stuck on this problem for like half an hour.

OpenStudy (amorfide):

good luck on your next question!

OpenStudy (anonymous):

Thanks